OpenStudy (seattle12345):

how do I solve this? Solve the rational equation: 9/(x-10)=5/(7x-5)

OpenStudy (mimi_x3):

\[\frac{9}{x-10} = \frac{5}{7x-5} \] Cross Multiply.

OpenStudy (mimi_x3):

\(9(7x-5)=5(x-10)\) Can you do it noow>

OpenStudy (anonymous):

x= -5/58

OpenStudy (maheshmeghwal9):